OpenStudy (anonymous):

Show geometrically why ∫sqrt(2-x^2) = (pi/4)+(1/2)

OpenStudy (anonymous):

\[\int\limits_{0}^{1}\sqrt{2-x ^{2}}dx = \Pi /4+1/2\]

OpenStudy (amistre64):

geometrically? wouldnt that just be a graph of the sqrt part?

OpenStudy (anonymous):

You have to like break it into two peices

OpenStudy (anonymous):

a sector of a circle and a triangle

OpenStudy (amistre64):

oh, then the graph would be useful at anyrate

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=sqrt%282-x%5E2%29%2C+real%2C+x%3D-2to2%2C+y%3D0to2 the wolf hates when I try to define stuff ....

OpenStudy (jamesj):

right exactly. The segment of the circle has angle pi/4, or 1/8 of a complete circle of 2pi and radius sqrt(2). Hence the segment of circle has area \[ \frac{1}{8}\pi r^2 = \frac{2}{8}\pi = \frac{\pi}{4} \] Now, what about the triangle?

OpenStudy (anonymous):

well the triangle wld be a special triangle

OpenStudy (jamesj):

No, it would be a right angled triangle.

OpenStudy (anonymous):

1.1 and sqrt(2)

OpenStudy (anonymous):

ya a right triangle

OpenStudy (jamesj):

what's the length of the base, b? And the height, h?

OpenStudy (anonymous):



OpenStudy (jamesj):

yes, exactly. And its area?

OpenStudy (anonymous):

(b*H)/2

OpenStudy (anonymous):

so it is a 1/2

OpenStudy (anonymous):

ohhhhh I see

OpenStudy (jamesj):

Nice question.

OpenStudy (anonymous):

Thanks :D